# The Structure of Hypergraphs Arising in Cellular Mobile Communication Systems

### v3

@@ 11 Hypergraphs Of Line Networks
In this section, we investigate line networks, which are wireless networks $\langle S,\alpha,\beta\rangle$ where the set $S$ of stations lies on a straight line such as the $x$-axis. Line networks are an interesting and simple special case to study first and have been studied by many researchers [35] [17]. We assume throughout that the path loss exponent $\alpha=4$.
@@
To motivate the problem studied in this section, first consider the simpler model of unit disk graphs. Let $S=\{s_{i}:i=1,\ldots,n\}$ be a set of stations, where the locations of the stations are given by $s_{i}=(i,0)(i=1,\ldots,n)$, so that the distance between any two consecutive stations is exactly $1$ unit. Under the unit disk graph model, the conflict graph would be the path graph $P_{n}$ on $n$ vertices. Hence, the path graph $P_{n}$ is a unit disk graph, and we would like to generalize this result. A natural question that arises is the following: Does the hypergraph $P_{n}$ arise as the interference model of a wireless network? In other words, is the path hypergraph $P_{n}$ realizable? If $S$ is assumed to be the set of $n$ locations above, then every pair of consecutive vertices is forbidden, and so $P_{n}$ is contained in the hypergraph generated by the wireless network $\langle S,\alpha,\beta\rangle$. The question is whether the hypergraph contains other minimal forbidden subsets besides the edge set of $P_{n}$. In other words, can the cumulative interference from multiple nodes, no two of which are consecutive, be intolerable? We answer this question in the negative, thereby proving that the hypergraph $P_{n}$ is realizable.
@@
Suppose the path loss exponent $\alpha=4$. Then, the path hypergraph $P_{n}$ is realizable for all $n\geq 1$.
Proof: A wireless network $\langle S,\alpha,\beta\rangle$ will be given that generates the path hypergraph $P_{n}$. The assertion clearly holds if $n=1$, so suppose $n\geq 2$. Without loss of generality, it can be assumed that the reception threshold $\beta=1$. Recall that $\alpha=4$. Let $S=\{s_{1},s_{2},\ldots,s_{n}\}$, where $s_{i}$ is at location $(i,0)$ for all $i$. Thus, the spacing between two consecutive stations on the line is exactly $1$. Let $H=(S,\mathcal{E})$ denote the hypergraph generated by the wireless network $\langle S,\alpha,\beta\rangle$. Because the distance between any two consecutive stations on the line is at most $1$ and the distance between any two nonconsecutive stations on the line is greater than $1$, the minimal forbidden sets of size $2$ are exactly the edges of the path graph $P_{n}$. In other words, $\mathcal{E}$ contains the edge set of $P_{n}$. It suffices to show that there are no other minimal forbidden sets.
@@
The first inequality above follows from the fact that for any station $s_{i}\in W$ whose location is $(i,0)$, the two stations in $W-\{s_{i}\}$ closest to $s_{i}$ would be at least as far apart from $s_{i}$ as $s_{i-2}$ and $s_{i+2}$ are, since the stations in $W$ are pairwise nonconsecutive. The next two stations in $W-\{s_{i}\}$ closest to $s_{i}$ would be $s_{i-4}$ and $s_{i+4}$ (or stations even further away from $s_{i}$ than these two stations). There are at most $n$ stations on either side of $s_{i}$. Thus, the interference at $s_{i}$ due to transmissions by the remaining stations in $W$ can be upper bounded as shown by the first inequality. The second inequality is clear because we are replacing the finite sum by an infinite series. The last equality follows from Euler’s formula that the sum of the reciprocals of the fourth powers of the positive integers is given by $\sum_{n=1}^{\infty}\frac{1}{n^{4}}=\frac{\pi^{4}}{90}$ [8] [9]. It follows that the hypergraph generated by the wireless network $\langle S,\alpha,\beta\rangle$ is the path hypergraph $P_{n}$.
Theorem 33, which states that the path hypergraph $P_{n}$ is realizable, assumed that the path loss exponent $\alpha=4$. If the path loss exponent $\alpha=2$, the assertion still holds. The same wireless network realization given in the proof of Theorem 33 can be used along with Euler’s formula $\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\frac{\pi^{2}}{6}$ to prove this assertion.
@@ 12 Concluding Remarks
Motivated by recent results that the worst-case performance of the distributed maximal scheduling algorithm under the hypergraph interference model is characterized by the interference degree of the hypergraph, in the present work, we investigated the properties of this hypergraph invariant, and in particular, the properties of hypergraphs that are realizable. The problem of computing the interference degree of a hypergraph was shown to be NP-hard and some properties of the interference degree of hypergraphs were proved.
It is known that the maximal value of $r$ such that the graph $K_{1,r}$ is a unit disk graph is $r=5$. An open problem in the literature is to extend the results in the literature on unit disk graphs to hypergraph models. In the present work, we investigated the additional structure possessed by hypergraphs that are realizable, i.e. that arise in practice due to physical constraints. The function $r(\alpha)$ was defined to be the maximal value of $k$ such that the $2$-uniform hypergraph $K_{1,k}$ is generated by some wireless network with path loss exponent $\alpha$. Motivated by previous results on the performance guarantees of distributed algorithms, in the present work we obtained the exact value of $r(\alpha)$ for various integer values of $\alpha$. These results were further extended to nonintegral values of the path loss exponent $\alpha$. We also investigated hypergraphs generated by line networks and showed that the path graph $P_{n}$ is realizable when $\alpha=4$.
There are several directions for further research. Corollary 32 gives the value of $r(\alpha)$ for certain intervals. Firstly, determining $r(\alpha)$ for the remaining values of the path loss exponent $\alpha$ and obtaining the transition values in Corollary 32 exactly using analytical methods are directions for further research. Secondly, determining which other structured hypergraphs besides the stars $K_{1,k}$ arise in practice as the interference model of wireless networks, based on physical constraints, is an open problem. Finally, it would be worth investigating the maximum possible interference degree of realizable hypergraphs, as this gives a performance guarantee on distributed algorithms for wireless networks under the hypergraph interference model.

### v4

@@ 11 Hypergraphs Of Line Networks
In this section, we investigate line networks, which are wireless networks $\langle S,\gamma,\beta\rangle$ where the set $S$ of stations lies on a straight line such as the $x$-axis. Line networks are an interesting and simple special case to study first and have been studied by many researchers [37] [17]. We assume throughout that the path loss exponent $\gamma=4$.
@@
To motivate the problem studied in this section, first consider the simpler model of unit disk graphs. Let $S=\{s_{i}:i=1,\ldots,n\}$ be a set of stations, where the locations of the stations are given by $s_{i}=(i,0)(i=1,\ldots,n)$, so that the distance between any two consecutive stations is exactly $1$ unit. Under the unit disk graph model, the conflict graph would be the path graph $P_{n}$ on $n$ vertices. Hence, the path graph $P_{n}$ is a unit disk graph, and we would like to generalize this result. A natural question that arises is the following: Does the hypergraph $P_{n}$ arise as the interference model of a wireless network? In other words, is the path hypergraph $P_{n}$ realizable? If $S$ is assumed to be the set of $n$ locations above, then every pair of consecutive vertices is forbidden, and so $P_{n}$ is contained in the hypergraph generated by the wireless network $\langle S,\gamma,\beta\rangle$. The question is whether the hypergraph contains other minimal forbidden subsets besides the edge set of $P_{n}$. In other words, can the cumulative interference from multiple nodes, no two of which are consecutive, be intolerable? We answer this question in the negative, thereby proving that the hypergraph $P_{n}$ is realizable.
@@
Suppose the path loss exponent $\gamma=4$. Then, the path hypergraph $P_{n}$ is realizable for all $n\geq 1$.
Proof: A wireless network $\langle S,\gamma,\beta\rangle$ will be given that generates the path hypergraph $P_{n}$. The assertion clearly holds if $n=1$, so suppose $n\geq 2$. Without loss of generality, it can be assumed that the reception threshold $\beta=1$. Recall that $\gamma=4$. Let $S=\{s_{1},s_{2},\ldots,s_{n}\}$, where $s_{i}$ is at location $(i,0)$ for all $i$. Thus, the spacing between two consecutive stations on the line is exactly $1$. Let $H=(S,\mathcal{E})$ denote the hypergraph generated by the wireless network $\langle S,\gamma,\beta\rangle$. Because the distance between any two consecutive stations on the line is at most $1$ and the distance between any two nonconsecutive stations on the line is greater than $1$, the minimal forbidden sets of size $2$ are exactly the edges of the path graph $P_{n}$. In other words, $\mathcal{E}$ contains the edge set of $P_{n}$. It suffices to show that there are no other minimal forbidden sets.
@@
The first inequality above follows from the fact that for any station $s_{i}\in W$ whose location is $(i,0)$, the two stations in $W-\{s_{i}\}$ closest to $s_{i}$ would be at least as far apart from $s_{i}$ as $s_{i-2}$ and $s_{i+2}$ are, since the stations in $W$ are pairwise nonconsecutive. The next two stations in $W-\{s_{i}\}$ closest to $s_{i}$ would be $s_{i-4}$ and $s_{i+4}$ (or stations even further away from $s_{i}$ than these two stations). There are at most $n$ stations on either side of $s_{i}$. Thus, the interference at $s_{i}$ due to transmissions by the remaining stations in $W$ can be upper bounded as shown by the first inequality. The second inequality is clear because we are replacing the finite sum by an infinite series. The last equality follows from Euler’s formula that the sum of the reciprocals of the fourth powers of the positive integers is given by $\sum_{n=1}^{\infty}\frac{1}{n^{4}}=\frac{\pi^{4}}{90}$ [8] [9]. It follows that the hypergraph generated by the wireless network $\langle S,\gamma,\beta\rangle$ is the path hypergraph $P_{n}$.
Theorem 36, which states that the path hypergraph $P_{n}$ is realizable, assumed that the path loss exponent $\gamma=4$. If the path loss exponent $\gamma=2$, the assertion still holds. The same wireless network realization given in the proof of Theorem 36 can be used along with Euler’s formula $\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\frac{\pi^{2}}{6}$ to prove this assertion.
@@ 12 Concluding Remarks
In the present work, we investigated the properties of a hypergraph invariant - the interference degree of a hypergraph. The problem of computing the interference degree of a hypergraph was shown to be NP-hard and some properties of the interference degree of hypergraphs were proved. Furthermore, we investigated the properties of hypergraphs that are realizable, i.e. that arise in practice due to physical (SINR) constraints. It is believed that the additional structure possessed by realizable hypergraphs can be exploited for efficient algorithm design.
An open problem in the literature is to extend the results in the literature on unit disk graphs to hypergraph models. It is known that the maximal value of $r$ such that the graph $K_{1,r}$ is a unit disk graph is $r=5$. In the present work, these results were extended to hypergraph models. More specifically, the function $r(\gamma)$ was defined to be the maximal value of $k$ such that the $2$-uniform hypergraph $K_{1,k}$ is generated by some wireless network with path loss exponent $\gamma$. We obtained the exact value of $r(\gamma)$ for various integer values of $\gamma$. These results were further extended to nonintegral values of the path loss exponent $\gamma$. We also investigated hypergraphs generated by line networks and showed that the path graph $P_{n}$ is realizable when $\gamma=4$.
There are several directions for further research. Corollary 35 gives the value of $r(\gamma)$ for certain intervals. Firstly, determining $r(\gamma)$ for the remaining values of the path loss exponent $\gamma$ and obtaining the transition values in Corollary 35 exactly using analytical methods are directions for further research. Secondly, determining which other structured hypergraphs besides the stars $K_{1,k}$ arise in practice as the interference model of wireless networks, based on physical constraints, is an open problem. Finally, it would be worth investigating the maximum possible interference degree of realizable hypergraphs, as this gives a performance guarantee on distributed algorithms for wireless networks under the hypergraph interference model.
13 Acknowledgements
Thanks are due to the anonymous reviewers for helpful comments and suggestions.